A) Consider the insoluble compound zinc hydroxide, Zn(OH)2. The zinc ion also forms a complex with ammonia. Write a balanced net ionic equation to show why the solubility of Zn(OH)2(s) increases in the presence of ammonia and calculate the equilibrium constant for this reaction.

For Zn(NH3)42+, Kf = 2.9×109. Specify states such as (aq) or (s) and provide K.

K = ______

B) Consider the insoluble compound nickel(II) hydroxide, Ni(OH)2. The nickel ion also forms a complex with cyanide ions. Write a balanced net ionic equation to show why the solubility of Ni(OH)2(s) increases in the presence of cyanide ions and calculate the equilibrium constant for this reaction.

For Ni(CN)42-, Kf = 1.0×1031. Specify states such as (aq) or (s) and provide K.

K = ______

C) Consider the insoluble compound zinc carbonate, ZnCO3. The zinc ion also forms a complex with hydroxide ions. Write a balanced net ionic equation to show why the solubility of ZnCO3(s) increases in the presence of hydroxide ions and calculate the equilibrium constant for this reaction.

For Zn(OH)42-, Kf = 2.9×1015. Specify states such as (aq) or (s) and provide K.

K = ______
Answer #1

A) Zn(OH)2 (s) Zn^2+ (aq) + 2OH^- (aq); Ksp = 4.5×10^(-17)

Zn^2+ (aq) + 4NH3 (aq) [Zn(NH3)4]^2+ (aq); Kf = 2.9×10^9

By adding the above two equations, you will get

Zn(OH)2 (s) + 4NH3(aq) [Zn(NH3)4]^2+ (aq) + 2OH^-; K = Ksp × Of = 1.305×10^(-7)

Similarly, you can approach for the remaining two parts B and C.